OpenStudy (anonymous):

Find the inverse, & determine if its a function. g(x)=(x-2)/2

OpenStudy (anonymous):

y=(x-2)/2 2y=x-2 2y+2=x x=2y+2

OpenStudy (anonymous):

i am letting g(x)=y

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

None of them are my answer choices tho,

OpenStudy (anonymous):

what are choices?

OpenStudy (anonymous):

2x+2 2/(x-2) 1/2x +1 -2x-2 x

OpenStudy (anonymous):

ya you have to interchange x and y in x=2y+2 which will give you y=2x+2

OpenStudy (anonymous):

oh yeah. Ok. Thanks.

OpenStudy (anonymous):

And it is a function right?

OpenStudy (anonymous):